# Learning Motion Blur Robust Vision Transformers with Dynamic Early Exit for Real-Time UAV Tracking

### v1

@@ Abstract
Recently, the surge in the adoption of single-stream architectures utilizing pre-trained ViT backbones represents a promising advancement in the field of generic visual tracking. By integrating feature extraction and fusion into a cohesive framework, these architectures offer improved performance, efficiency, and robustness. However, there has been limited exploration into optimizing these frameworks for UAV tracking. In this paper, we boost the efficiency of this framework by tailoring it into an adaptive computation framework that dynamically exits Transformer blocks for real-time UAV tracking. The motivation behind this is that tracking tasks with fewer challenges can be adequately addressed using low-level feature representations. Simpler tasks can often be handled with less demanding, lower-level features. This approach allows the model use computational resources more efficiently by focusing on complex tasks and conserving resources for easier ones. Another significant enhancement introduced in this paper is the improved effectiveness of ViTs in handling motion blur, a common issue in UAV tracking caused by the fast movements of either the UAV, the tracked objects, or both. This is achieved by acquiring motion blur robust representations through enforcing invariance in the feature representation of the target with respect to simulated motion blur. The proposed approach is dubbed BDTrack. Extensive experiments conducted on five tracking benchmarks validate the effectiveness and versatility of our approach, establishing it as a cutting-edge solution in real-time UAV tracking. Code is released at: https://github.com/wuyou3474/BDTrack.
Index Terms:
@@
UAV tracking, real-time, Vision Transformer, dynamic early exiting, motion blur robust representations.
I Introduction
@@
Unmanned Aerial Vehicle (UAV) tracking involves the process of determining and predicting the location and size of a specific object in successive aerial images. This capability has become increasingly essential in a variety of fields, such as public safety, disaster management, environmental monitoring, industrial inspections, and agriculture, due to its wide-ranging applications and benefits [1, 49, 18, 4, 12, 2]. Despite its benefits, UAV tracking faces significant challenges, including extreme viewing angles that require robust algorithms to handle diverse perspectives, severe motion blur from rapid UAV movement that complicates the tracking process, and significant occlusions where objects are obscured by other elements. Additionally, UAVs have limited computational resources and battery life, necessitating highly efficient tracking algorithms to ensure sustained and accurate operation during missions [5, 6]. Therefore, a practical tracker must ensure accurate tracking while addressing the computational and power constraints of the UAVs. Although discriminative correlation filters (DCF)-based trackers are widely used in UAV tracking for their efficiency [7, 9, 11, 3], their tracking precision often lags behind that of deep learning (DL)-based methods [49, 5, 4, 12]. A recent trend in the generic tracking community sees a growing preference for single-stream architectures, seamlessly integrating feature extraction and correlation using pre-trained Vision Transformer (ViT) backbone networks [13, 14, 17, 8], showcasing notable success. In line with this trend, Aba-ViTrack [18] presents an efficient DL-based tracker that utilizes this framework. It includes an adaptive and background-aware token computation method to reduce inference time, demonstrating impressive precision and speed, making it a state-of-the-art (SOTA) tracker for real-time UAV tracking. However, the varying token number of this method entails unstructured access operations, resulting in a notable time costs. In our research, we employ the same single-stream architecture, but our focus is on enhancing the efficiency of ViTs through more structured methods: the early exiting strategy.
@@
Figure 1: Compared with state-of-the-art UAV tracking algorithms on UAVDT. Our BDTrack-DeiT sets a new record with 0.841 precision and still runs efficiently at around 305 FPS.
@@
Early exiting is a technique designed to expedite the processing of input data by allowing a model to terminate its forward pass prematurely based on certain criteria, reducing computational costs without compromising much predictive accuracy, especially if the behavior or characteristics of the task vary based on the specific examples or instances involved. This approach has gained traction in various domains, especially in real-time applications and resource-constrained environments, and a number of studies have explored its implications [19, 20, 21, 40, 22]. Research in this area explores various strategies, including adaptive methods that dynamically assess model confidence [19, 22], distillation [23], architecture design [24, 21], and training scheme [20, 40]. This approach has been applied across various deep learning architectures, including convolutional neural networks (CNNs) [19, 23], recurrent neural networks (RNNs) [24], and Transformer-based models [21, 22]. While it has found success in applications ranging from natural language processing [24, 40] to many vision tasks (e.g., image classification [22], video recognition [25], and image captioning [26]) and vision language learning [27], its potential for enhancing efficiency in UAV tracking has not been explored to the best of our knowledge. In this work, we introduce a Dynamic Early Exit Module (DEEM) to streamline the architecture of ViTs and enhance the efficiency of the inference process for real-time UAV tracking. Given that the tracking challenge for a specific target is context-dependent, the adaptive neglect of high-level ViT blocks can significantly reduce unnecessary high-level semantic features without significantly affecting the performance of the tracker.
@@
In addition, this research also aims to enhance the performance of UAV trackers by mitigating the adverse effects of motion blur, which is a prevalent challenge in UAV tracking as rapid and unpredictable movements of UAVs are common due to factors such as wind, changes in altitude, sudden maneuvers, or the need to track fast-moving targets [11, 6, 28]. Although numerous methods aiming to address motion blur in visual tracking have been proposed [29, 30, 31, 32, 33], their broad applicability is impeded by significant drawbacks. Their intricate designs, complex architectures [34, 33], or training procedures [31, 33], present implementation challenges and requires substantial computational resources. Moreover, many of them neglect efficiency considerations [29, 30, 32], rendering them less suitable for resource-constrained scenarios such as real-time UAV tracking. Simplifying the design and ensuring efficiency are pivotal for increasing the adoption and effectiveness of methods designed to tackle motion blur in UAV tracking, especially for real-time needs. In this work, we propose to learn motion blur robust feature representations with ViTs by minimizing the mean square error (MSE) between the feature representation of the original template and its blurred version subject to motion blur. Our approach distinguishes itself through its simplicity, avoiding the inclusion of extra architectural complexities. Instead, it introduces a straightforward yet effective enhancement by incorporating only an additional loss. This enhancement can be seamlessly integrated into other tracking methods to bolster robustness against motion blur, which imposes no extra computational burden during the inference phase, making it particularly well-suited for real-time UAV tracking scenarios where efficiency is crucial. It is worth noting that the targets subjecting to motion blur is simulated considering that this manner can be conducted on-demand and provides a cost-effective and scalable solution, among other advantages [35]. The proposed framework is called BDTrack. Validation using real-world data demonstrates that our method significantly enhances the robustness against motion blur of baseline methods. Remarkably, our BDTrack-DeiT outperforms the SOTA tracker ABDNet [33], which was proposed to address motion blur, by 5.5% in precision on the motion blur subset of the UAVDT dataset with more than 2 times faster GPU speed (See Fig. 4 and Table I). Extensive experiments on five benchmarks show that our method achieves SOTA performance. As shown in Fig. 1, our method sets a new record with a precision of 0.841 and runs efficiently at around 305 frames per second (FPS) on the UAVDT. The following is a summary of our contributions:
@@
  • We introduce a novel module that adaptively streamlines the architecture of Vision Transformers (ViTs) by dynamically exiting Transformer blocks. Our method is straightforward and can significantly enhance the efficiency of baseline ViT-based UAV tracking methods with minimal impact on their tracking performance.
@@
  • We make the first attempt to learn motion blur robust ViTs by enforcing invariance in the feature representation of the target against simulated motion blur using simply a mean square error (MSE) loss. Our approach incurs no additional computational burden during the inference, making it well-suited for real-time UAV tracking.
@@
  • We present BDTrack, an efficient tracker seamlessly integrating these components, which is readily integrable into similar ViT-based trackers. BDTrack is able to maintain high tracking speeds while delivering exceptional performance. Extensive experiments on five benchmarks confirm its state-of-the-art performance.
The remainder of this paper is structured as follows: Section II provides an overview of previous research related to this study. In Section III, we describe the methodology of the proposed technique in detail. Section IV outlines the experiments conducted and discusses their results. Finally, Section V presents the conclusions and insights drawn from the paper.
II Related Work
II-A Motion Blur Aware Trackers
@@
Motion blur presents a significant challenge in visual tracking, and various approaches have been proposed to address this issue and enhance the robustness of tracking algorithms [10, 30, 31, 32, 33], including employing specialized architectures or mechanisms to handle motion blur, learning motion blur robust representations, and combining motion deblurring techniques. For example, Ding et al. [10] introduce a method to enhance the robustness of the appearance model against severe blur by incorporating various blur kernels during the training process. Guo et al. [32] propose a novel Generative Adversarial Network (GAN)-based scheme aimed at enhancing tracker robustness to motion blurs. To address the challenges posed by motion blur in UAV tracking, Zuo et al. [33] introduced the Tracking-Oriented Adversarial Blur-Deblur Network (ABDNet) very recently, which includes a novel deblurring component to restore the visual appearance of blurred targets and an innovative blur generator that creates realistic blurry images for adversarial training. While existing methods have shown success in addressing motion blur in visual tracking, their implementation is complex and the efficiency of the algorithms is not a primary consideration. This complexity and inefficiency make these methods unsuitable for real-time UAV tracking scenarios, where computational resources and time are limited. In this study, we introduce motion blur robust ViTs for UAV tracking by learning motion blur invariant feature representations in a simpler and more efficient manner. Our approach is designed to incur no additional computational burden during the inference phase, making it exceptionally well-suited for real-time UAV tracking scenarios where efficiency is paramount. Furthermore, our method can be easily adapted to other ViT-based tracking methods, enhancing their performance in the presence of motion blur without compromising speed or computational efficiency.
II-B Efficient Vision Transformers
@@
With the increased popularity, efficient ViTs based on conditional computation have very recently explored adaptive inference for model acceleration, which dynamically adjust the computational load based on input complexity, enabling ViTs to allocate resources judiciously during inference. For instance, DynamicViT [38] introduces extra control gates to pause specific tokens using the Gumbel-softmax trick. A-ViT [39] adopts an Adaptive Computation Time (ACT)-like approach to eliminate the need for extra halting sub-networks, resulting in simultaneous improvements in efficiency, accuracy, and token importance allocation. A-ViT was also exploited in Aba-ViTrack [18] to built efficient trackers for real-time UAV tracking. However, its variable number of tokens incurs considerable time cost due to additional unstructured assess operations. More structured methods can resort to early exiting that leverage dynamic early exiting strategies to improve inference efficiency of Transformer models, as demonstrated in [40, 26, 27, 22]. Nevertheless, these studies primarily focus on large-scale models and incorporate inefficient architectures or training methods for making early exiting decisions. These criteria for early exiting are challenging to adapt effectively to lightweight models. For example, the early exiting strategy proposed for vision language models in [27] relies on layer-wise Cosine similarities. While Cosine similarities have proven effective for large-scale models characterized by high-level and high-dimensional feature representations, their performance may be suboptimal in lightweight models where non-linearities and complex dependencies are challenging to linearize or trivialize [41, 42]. The dynamic early exiting method proposed for accelerating ViTs in [22] involves training delicate Local Perception and Global Aggregation Heads separately with distinct supervisions. However, this intricate training setup introduces significant inefficiencies into the training process. In this work, we explore a more structured and efficient approach to dynamically early exit ViT blocks by incorporating a dynamic early exit module. Our method aims to simplify the training process, improve inference efficiency, and maintain adaptability across different ViT-based models.
Figure 2: Overview of our BDTrack. It comprises of a single-stream backbone and a prediction head. Simulated blurred templates and Dynamic Early Exit Modules (see Fig. 3) are employed to learn motion blur robust representations and to boost efficiency, respectively.
III Method
@@ III-A Overview
III-B Learning Motion Blur Robust ViTs (MBRV)
@@
A motion blur robust ViT (MBRV) serves as a pivotal component in our pursuit of a compact and efficient end-to-end tracker specifically designed for real-time UAV tracking. In the context of UAV tracking, where rapid and accurate object localization is paramount, the presence of motion blur in captured images poses a significant challenge. Motion blur can distort object features, leading to inaccuracies in tracking algorithms and potentially compromising the effectiveness of the entire tracking system. The inclusion of MBRV addresses these challenges by ensuring that the ViT-based backbone of our tracking system remains effective even in the presence of motion blur. By minimizing the Mean Squared Error (MSE) between the feature representations of the target template and its blurred counterpart, MBRV enhances the robustness of the tracker. Furthermore, the compact and efficient nature of MBRV aligns with our goal of developing a tracker optimized for real-time UAV tracking.
@@
During the training phase, the MBRV receives three inputs: a target template $Z$, a blurred target template $Z^{\prime}$= $\mathfrak{m}(Z)$, and a search image $X$, where $\mathfrak{m}(Z)$ represents applying simulated motion blur to the template image $Z$ using the linear motion blur method implemented in [43]. The input images are first segmented and then flattened into sequences of patches. Subsequently, the sequences corresponding to $Z$ and $X$ are concatenated into an ordered sequence denoted by $[Z,X]$. The same procedure is applied to the sequences corresponding to $Z^{\prime}$ and $X$, leading to $[Z^{\prime},X]$. In the forward pass, $[Z,X]$ is tokenized by a trainable linear projection layer $\mathfrak{E}$ called patch embedding, which generates $\mathcal{K}$ tokens as follows:
$\textbf{t}_{1:\mathcal{K}}^{0}=\mathfrak{E}([Z,X])\in\mathbb{R}^{\mathcal{K}\times d},$ (1)
@@
where the embedding dimension of each token is denoted by $d$, and the token sequences $t^{0}_{1:\mathcal{K}z}$ and $t^{0}_{\mathcal{K}_{z}+1:\mathcal{K}}$ ($1<\mathcal{K}_{z}<\mathcal{K}_{x}<\mathcal{K}$) correspond to the template and search image, respectively, where $\mathcal{K}=\mathcal{K}_{z}+\mathcal{K}_{x}$. The $\mathcal{K}$ tokens $\textbf{t}_{1:\mathcal{K}}^{0}$ are then input into the Transformer blocks to obtain their final feature representations. Let $\mathfrak{F}^{l}$ represent the Transformer block at layer $l$, transforming all tokens from layer $(l-1)$ as $\textbf{t}_{1:\mathcal{K}}^{l}=\mathfrak{F}^{l}(\textbf{t}_{1:\mathcal{K}}^{l-1})$. The overall MBRV of $L$ ViT blocks in total, denoted by $\mathfrak{B}$, can be expressed as:
$\textbf{t}_{1:\mathcal{K}}^{L}(Z,X)=\mathfrak{B}(Z,X)=\mathfrak{F}^{L}\circ\mathfrak{F}^{L-1}\circ...\circ\mathfrak{F}^{1}\circ\mathfrak{E}([Z,X]),$ (2)
@@
where $\circ$ denotes the composition operation. Similarly, we can derive the feature representation of $[Z^{\prime},X]$. As $[Z,X]$ and $[Z^{\prime},X]$ are ordered sequences of tokens, the feature representation of $Z$ and $Z^{\prime}$ can be retrieved by tracking their token indices in the respective ordered sequences, specifically denoted as $\textbf{t}_{1:\mathcal{K}z}^{L}(Z,X)$ and $\textbf{t}_{1:\mathcal{K}_{z}}^{L}(Z^{\prime},X)$, respectively. The core idea behind MBRV is to minimize the mean square error between the feature representation of $Z$ and that of $Z^{\prime}$. This is achieved by minimizing the following MSE loss:
$\mathcal{L}_{br}=||\textbf{t}_{1:\mathcal{K}_{z}}^{L}(Z,X)-\textbf{t}_{1:\mathcal{K}_{z}}^{L}(Z^{\prime},X)||^{2}.$ (3)
@@
Figure 3: The detailed structure of the Dynamic Early Exit Module.
III-C Dynamic Early Exit Module (DEEM)

### v2

# Learning Motion Blur Robust Vision Transformers for Real-Time UAV Tracking

@@ Abstract
Unmanned aerial vehicle (UAV) tracking is critical for applications like surveillance, search-and-rescue, and autonomous navigation. However, the high-speed movement of UAVs and targets introduces unique challenges, including real-time processing demands and severe motion blur, which degrade the performance of existing generic trackers. While single-stream vision transformer (ViT) architectures have shown promise in visual tracking, their computational inefficiency and lack of UAV-specific optimizations limit their practicality in this domain. In this paper, we boost the efficiency of this framework by tailoring it into an adaptive computation framework that dynamically exits Transformer blocks for real-time UAV tracking. The motivation behind this is that tracking tasks with fewer challenges can be adequately addressed using low-level feature representations. Simpler tasks can often be handled with less demanding, lower-level features. This approach allows the model use computational resources more efficiently by focusing on complex tasks and conserving resources for easier ones. Another significant enhancement introduced in this paper is the improved effectiveness of ViTs in handling motion blur, a common issue in UAV tracking caused by the fast movements of either the UAV, the tracked objects, or both. This is achieved by acquiring motion blur robust representations through enforcing invariance in the feature representation of the target with respect to simulated motion blur. We refer to our proposed approach as BDTrack. Extensive experiments conducted on four tracking benchmarks validate the effectiveness and versatility of our approach, demonstrating its potential as a practical and effective approach for real-time UAV tracking. Code is released at: https://github.com/wuyou3474/BDTrack.
keywords:
@@
UAV tracking , real-time , Vision Transformer , dynamic early exiting , motion blur robust representations
journal: Expert Systems With Applications
[…]
1 Introduction
@@
In recent years, Artificial Intelligence (AI) has brought transformative progress across various domains, including AI-driven models [1, 2], multimodal learning [3, 4], dynamic resource allocation [5, 6], unmanned aerial vehicle (UAV) tracking [7, 8, 9], and intelligent energy systems and control applications [10, 11, 12, 13, 14, 15, 16]. Recent advancements have broadly improved system performance, energy efficiency, and real-time decision-making across various engineering fields. These developments have motivated us to apply them to UAV tracking, aiming to improve accuracy, enhance robustness in dynamic environments, and achieve efficient real-time processing. UAV tracking involves the process of determining and predicting the location and size of a specific object in successive aerial images. This capability has become increasingly essential in a variety of fields, such as public safety, environmental monitoring, industrial inspections, disaster relief, and agriculture, due to its wide-ranging applications and benefits [7, 17, 9, 18, 19, 20]. These applications present unique challenges, including the need for accurate tracking in resource-constrained environments, where UAVs often face limited battery life and computational resources. Therefore, efficient tracking algorithms that ensure sustained and accurate performance under these conditions are paramount [8, 21]. Traditional discriminative correlation filter (DCF)-based trackers have been widely used in UAV tracking due to their efficiency [22, 23, 24, 25]. However, their tracking precision often falls behind that of deep learning (DL)-based methods, which have shown promising results in terms of precision [17, 8, 18, 19]. Despite the advantages of DL-based methods, they face significant challenges in real-time applications, especially in resource-limited scenarios, where both computational power and battery life are constrained. Recent advances in the tracking community have led to the emergence of single-stream architectures, particularly those leveraging pre-trained vision transformer (ViT) backbone networks. These architectures have demonstrated notable success in integrating feature extraction and correlation [26, 27, 28, 29]. In line with this trend, Aba-ViTrack [9] introduces an efficient DL-based tracker that adopts a single-stream ViT framework with adaptive token selection, achieving a good balance between precision and speed. However, its variable token scheme incurs unstructured access overhead. Motivated by this, we adopt a similar architecture but enhance efficiency through a structured early exiting strategy.
@@
Figure 1: Compared with state-of-the-art UAV tracking algorithms on four UAV tracking benchmarks, our BDTrack-DeiT sets a new record with 0.844 precision while maintaining efficient performance at around 283 FPS.
@@
Early exiting is a technique designed to expedite the processing of input data by allowing a model to terminate its forward pass prematurely based on certain criteria, reducing computational costs without compromising much predictive accuracy, especially if the behavior or characteristics of the task vary based on the specific examples or instances involved. While previous research has mainly explored this idea in domains such as natural language processing and image classification [30, 31, 32, 33, 6, 34], visual tracking, particularly real-time UAV tracking, poses unique challenges that differentiate it from these static tasks. Unlike image classification, visual tracking requires maintaining temporal consistency across frames, coping with abrupt target motion, deformation, scale variation, and appearance changes. These properties demand early-exit strategies that not only assess the model’s confidence for a single frame but also account for the temporal dependency of predictions and the evolving target appearance over time. Existing studies have investigated diverse strategies, including adaptive confidence-based exits [30, 6], distillation [35], architectural design [36, 32], and specialized training schemes [31, 33], applied to architectures ranging from convolutional neural networks (CNNs) [30, 35] to recurrent neural networks (RNNs) [36] and Transformers [32, 6]. While these methods have shown success in applications such as NLP [36, 33], image classification [6], video recognition [37], image captioning [38], and vision-language learning [5], their direct application to UAV tracking remains largely unexplored. This gap motivates our work, which adapts early-exit mechanisms to the real-time, temporally dependent, and appearance-variant nature of UAV tracking.
@@
In addition, this research also aims to enhance the performance of UAV trackers by mitigating the adverse effects of motion blur, which is a prevalent challenge in UAV tracking as rapid and unpredictable movements of UAVs are common due to factors such as wind, changes in altitude, sudden maneuvers, or the need to track fast-moving targets [24, 21, 39]. For example, existing trackers  [17, 8] show a marked increase in center location error (CLE) during real-world testing with motion-blurred sequences, rising from under 10 pixels to over 20, a threshold commonly regarded as tracking failure. This degradation becomes especially severe during aggressive UAV maneuvers such as sharp turns or rapid accelerations, where motion blur is significantly intensified, further impairing tracking performance. Although numerous methods have been proposed to address motion blur in visual tracking [40, 41, 42, 43, 44], their practical use in real-time UAV tracking remains limited due to inherent shortcomings. In contrast, we propose to learn motion blur robust feature representations with ViTs by minimizing the mean square error (MSE) between the feature representation of the original template and its blurred version subject to motion blur. Our approach distinguishes itself through its simplicity, avoiding the inclusion of extra architectural complexities. The proposed framework is called BDTrack. Validation using real-world data demonstrates that our method significantly enhances the robustness against motion blur of baseline methods. Remarkably, our BDTrack-DeiT outperforms the SOTA tracker ABDNet [44], which was proposed to address motion blur, by 5.5% in precision on the motion blur subset of the UAVDT dataset with more than 2 times faster GPU speed (See Fig. 4 and Table 1). Extensive experiments on four benchmarks show that our method achieves SOTA performance. As shown in Fig. 1, our method sets a new record with a average precision of 0.844 and runs efficiently at around 283 frames per second (FPS) on the four UAV tracking benchmarks. Its real-time, motion-blur-robust tracking improves UAV efficiency in applications such as agricultural inspection by enabling faster and wider-area coverage with minimal computational load, and in disaster relief operations by maintaining reliable tracking in challenging conditions to support quicker and more effective search-and-rescue missions. We also emphasize its suitability for resource-constrained UAVs through adaptive computation that preserves battery life and processing power.
The following is a summary of our contributions:
  • 1.
@@
    We introduce a novel module that adaptively streamlines the architecture of ViTs by dynamically exiting Transformer blocks. Our method is straightforward and can significantly enhance the efficiency of baseline ViT-based UAV tracking methods with minimal impact on their tracking performance.
  • 2.
@@
    We make a pioneering effort to enhance the motion blur robustness of vision transformers (ViTs) for real-time UAV tracking by leveraging a simple mean square error (MSE) loss to enforce feature invariance of the target under simulated blur. Our approach incurs no additional computational burden during the inference, making it well-suited for real-time UAV tracking.
  • 3.
@@
    We present BDTrack, an efficient tracker seamlessly integrating these components, which is readily integrable into similar ViT-based trackers. BDTrack is able to maintain high tracking speeds while delivering exceptional performance. Extensive experiments on four benchmarks confirm its state-of-the-art performance.
2 Related Work
2.1 Motion Blur Aware Trackers
@@
Motion blur presents a significant challenge in visual tracking, and various approaches have been proposed to address this issue and enhance the robustness of tracking algorithms [45, 41, 42, 43, 44], including employing specialized architectures or mechanisms to handle motion blur, learning motion blur robust representations, and combining motion deblurring techniques. For example, Ding et al. [45] introduce a method to enhance the robustness of the appearance model against severe blur by incorporating various blur kernels during the training process. Guo et al. [43] propose a novel generative adversarial network (GAN)-based scheme aimed at enhancing tracker robustness to motion blurs. To address the challenges posed by motion blur in UAV tracking, Zuo et al. [44] introduced the tracking-oriented adversarial blur-deblur network (ABDNet) very recently, which includes a novel deblurring component to restore the visual appearance of blurred targets and an innovative blur generator that creates realistic blurry images for adversarial training. However, existing methods for handling motion blur often involve complex architectures [46, 44] or training pipelines [42, 44] and neglect runtime efficiency [40, 41, 43],limiting their suitability for UAV applications. We introduce motion blur robust ViTs for UAV tracking that learn invariant feature representations efficiently without adding inference overhead, making them ideal for real-time scenarios. The method is easily adaptable to other ViT-based trackers, improving performance under motion blur without sacrificing speed.
2.2 Efficient Vision Transformers
@@
Conditional computation-based ViTs have recently enabled adaptive inference by dynamically adjusting computation based on input complexity [50, 51]. For example, A-ViT [51] eliminates extra halting networks via ACT-like mechanisms, improving both efficiency and accuracy. Aba-ViTrack [9] extends this design to UAV tracking, but suffers from the overhead of unstructured token selection. In contrast, our DEEM adopts a structured early-exit strategy that enables efficient inference, implemented using only a simple linear layer with sigmoid activation. Nevertheless, existing studies primarily focus on large-scale models and incorporate inefficient architectures or training methods for making early exiting decisions [33, 38, 5, 6]. These criteria for early exiting are challenging to adapt effectively to lightweight models. For example, the early exiting strategy proposed for vision language models in [5] relies on layer-wise cosine similarities. While cosine similarities have proven effective for large-scale models characterized by high-level and high-dimensional feature representations, their performance may be suboptimal in lightweight models where non-linearities and complex dependencies are challenging to linearize or trivialize [52, 53]. The dynamic early exiting method proposed for accelerating ViTs in [6] involves training delicate Local Perception and Global Aggregation Heads separately with distinct supervisions. However, this intricate training setup introduces significant inefficiencies into the training process. In this work, we explore a more structured and efficient approach to dynamically early exit ViT blocks by incorporating a dynamic early exit module. Our method aims to simplify the training process, improve inference efficiency, and maintain adaptability across different ViT-based models.
Figure 2: Overview of the proposed BDTrack. It comprises of a single-stream backbone and a prediction head. MBRV and DEEM (see Fig. 3) are employed to learn motion blur robust representations and to boost efficiency, respectively. Note that $\mathfrak{m}$ applies simulated motion blur to the template image $Z$ using the method from [54]. DEEM takes a subset of tokens from the target template and search image as input, outputting the exit score (i.e., $q$) at the corresponding ViT block.
3 Method
@@ 3.1 Overview
3.2 Learning Motion Blur Robust ViTs (MBRV)
@@
A motion blur robust ViT (MBRV) serves as a pivotal component in our pursuit of a compact and efficient end-to-end tracker specifically designed for real-time UAV tracking. In the context of UAV tracking, where rapid and accurate object localization is paramount, the presence of motion blur in captured images poses a significant challenge. Motion blur can distort object features, leading to inaccuracies in tracking algorithms and potentially compromising the effectiveness of the entire tracking system. The inclusion of MBRV addresses these challenges by ensuring that the ViT-based backbone of our tracking system remains effective even in the presence of motion blur. By minimizing the Mean Squared Error (MSE) between the feature representations of the target template and its blurred counterpart, MBRV enhances the robustness of the tracker. Furthermore, the compact and efficient nature of MBRV aligns with our goal of developing a tracker optimized for real-time UAV tracking. It is worthy of note that we choose to blur the template patch rather than the search patch is based on the following considerations: 1) It allows a controllable blurring of the target, as the object in the template always is sharp and unblurred while the object in the search patch is frequently subject to motion blur. 2) Since the size of the template patch and the scale of the target are fixed, the tokens associated with the template can be precisely identified. But the identification of the objects in the search patch may involve interpolation due to translation and scale variations of the objects. As the size of the search patch is fixed, interpolation has to be conducted on the feature space. However, interpolating in the feature space can lead to several issues, including loss of spatial information, misalignment of features, and inconsistent semantic meaning. These factors can result in poorer quality interpolated features, affect training dynamics, and potentially lead to suboptimal model performance.
@@
During the training phase, the MBRV receives three inputs: a target template $Z$, a blurred target template $Z^{\prime}$= $\mathfrak{m}(Z)$, and a search image $X$, where $\mathfrak{m}(Z)$ represents applying simulated motion blur to the template image $Z$ using the linear motion blur method implemented in [54]. The input images are first segmented and then flattened into sequences of patches. Subsequently, the sequences corresponding to $Z$ and $X$ are concatenated into an ordered sequence denoted by $[Z,X]$. The same procedure is applied to the sequences corresponding to $Z^{\prime}$ and $X$, leading to $Z^{\prime},X]$. In the forward pass, $[Z,X]$ is tokenized by a trainable linear projection layer $\mathcal{E}$ called patch embedding, which generates $\mathcal{K}$ tokens as follows:
$\textbf{t}_{1:\mathcal{K}}^{0}=\mathcal{E}([Z,X])\in\mathbb{R}^{\mathcal{K}\times d},$ (1)
@@
where the embedding dimension of each token is denoted by $d$, and the token sequences $t^{0}_{1:\mathcal{K}z}$ and $t^{0}_{\mathcal{K}_{z}+1:\mathcal{K}}$ ($1<\mathcal{K}_{z}<\mathcal{K}_{x}<\mathcal{K}$) correspond to the template and search image, respectively, where $\mathcal{K}=\mathcal{K}_{z}+\mathcal{K}_{x}$. The $\mathcal{K}$ tokens $\textbf{t}_{1:\mathcal{K}}^{0}$ are then input into the Transformer blocks to obtain their final feature representations. Let $\mathfrak{T}^{l}$ represent the Transformer block at layer $l$, transforming all tokens from layer (l1)(l-1) as $\textbf{t}_{1:\mathcal{K}}^{l}=\mathfrak{T}^{l}(\textbf{t}_{1:\mathcal{K}}^{l-1})$. The overall MBRV of $L$ ViT blocks in total, denoted by $\mathfrak{B}$, can be expressed as:
$\textbf{t}_{1:\mathcal{K}}^{L}(Z,X)=\mathfrak{B}(Z,X)=\mathfrak{T}^{L}\circ\mathfrak{T}^{L-1}\circ...\circ\mathfrak{T}^{1}\circ\mathcal{E}([Z,X]),$ (2)
@@
where $\circ$ denotes the composition operation. Similarly, we can derive the feature representation of $Z^{\prime},X]$. As $[Z,X]$ and $Z^{\prime},X]$ are ordered sequences of tokens, the feature representation of $Z$ and $Z^{\prime}$ can be retrieved by tracking their token indices in the respective ordered sequences, specifically denoted as $\textbf{t}_{1:\mathcal{K}z}^{L}(Z,X)$ and $\textbf{t}_{1:\mathcal{K}_{z}}^{L}(Z^{\prime},X)$, respectively. The core idea behind MBRV is to minimize the mean square error between the feature representation of $Z$ and that of $Z^{\prime}$. This is achieved by minimizing the following MSE loss:
[…]
@@
Figure 3: The detailed structure of the DEEM. The variables $\textbf{e}_{1}^{\textup{T}}$ represents a subset of all tokens output by the (i1)(i-1)-th ViT block, $\mathfrak{e}_{i}$ denotes the exit score of the $i$-th Transformer block, Sigmoid represents the sigmoid function, and $q^{i}$ denotes the cumulative exit score before (i1)(i-1).
3.3 Dynamic Early Exit Module (DEEM)